OpenStudy (anonymous):

Help will Give Medal △ABC has vertices of A (0, 0), B (4, 4), and C (8, 0). Find the orthocenter of △ABC. A.(0, 0) B.;(4, 4) C.(6, 2) D.(3, 1)

OpenStudy (anonymous):

Can one of you help me?

OpenStudy (anonymous):



OpenStudy (anonymous):

i cant see your drawing :/

OpenStudy (anonymous):

can anyone else help?

OpenStudy (anonymous):

I'll show examples, You can solve. OK. :))) We have given a triangle ABC whose vertices are(0, 6),(4, 6), (1, 3) In Step 1 we find slopes Of AB, BC,CA Slope formulae y2-y1⁄ x2-X1 slope AB= 6-6/4-0 = 0/4 =0 .... BC= 3-6/ 1-4 = -3/-3 =1 ....... CA=6-3/ 0-1 =3/-1 =-3 In Step 2 But we know Orthocentre is the point where perpendeculars drawn from vertex to opposite side meet. So Let's think a triangle ABC and AD, BE, CF are perpendiculars drawn to the vertex. Slope AD = -1/slope BC = -1/1 =-1 .......BE = -1/slope CA = -1/-3 = 1/3 .....CF = -1/slope AB = -1/0 undefined Step 3 we have now vertices and slopes of AD,BE,CF we find equations of lines AD,BE and CF we have A(0,6) and m =-1 we substitute in the equation y-y1 = m(x-X1) y-6=-1(x-0) y+x=6 - eq 1 B(4,6) and slope BE (1/3) y-6=1/3(x-4) 3Y-18=x-4 3y-x=14 -eq 2 C(1,3) and whose slope CF undefined So line is vertical and x=1 is the eq Now solving any of equations 1&2 we get values for( x,y) orhto centre (x,y) =( solving eq 1 and eq 2 we get x=1, y=5 ( I hope this helps you

OpenStudy (anonymous):

as an example.. :))